### v1

# The Echo 12–23 Texture: A Novel Flavour Paradigm for Neutrinos

Abstract
We construct a flavour guided model that realises the distinctive Echo 12–23 Texture in the neutrino mass matrix through a non-trivial interplay of symmetries. While the model accounts for charged lepton mass hierarchy, the texture offers interesting insights specifically into neutrino mass ordering and flavour dynamics. With clear imprints on low energy observables, the framework provides a minimal yet testable path toward understanding the origin of neutrino properties.
I Introduction
The Standard Model (SM) [1, 2, 3, 4, 5] of particle physics, based on the gauge group $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$, has successfully described a wide range of phenomena involving fundamental particles and interactions. However, several observations remain unexplained within its framework, including the origin of neutrino [6] masses, the nature of their mass ordering, and the hierarchy among charged lepton masses. Notably, neutrino oscillation experiments have firmly established that neutrinos are massive and undergo flavour transitions [7, 8, 9], necessitating an extension beyond the SM. Among various beyond SM (BSM) approaches, seesaw mechanisms [10, 11] provide elegant explanations for the smallness of neutrino masses. In particular, the Type-I seesaw [12, 13] introduces heavy right handed neutrinos ($\nu_{R}$), while the Type-II seesaw [13, 14] relies on an $SU(2)_{L}$ scalar triplet ($\Delta$) acquiring a small vacuum expectation value (vev). The typical Feynman diagrams for the Type-I and Type-II seesaw are shown in Figure. 1. An interesting scenario combining both, commonly known as the Type-I+II seesaw mechanism [15, 16, 17, 18], offers increased structural richness and emerges naturally in symmetry driven frameworks. The neutrino mass matrix, $M_{\nu}$, encodes valuable information about mixing angles ($\theta_{12},\theta_{13},\theta_{23}$), CP violating phases ($\delta,\alpha,\beta$), and the neutrino mass eigenvalues ($m_{1},m_{2},m_{3}$). However, neutrino oscillation data constrains only a subset of these parameters. To bridge this gap, various theoretical strategies have been explored, most notably the introduction of textures, where specific correlations or patterns in the matrix elements reduce parameter freedom and enhances the predictive power [19, 20, 21, 22, 23].
In this work, we propose a flavour motivated model embedded within a Type-I+II seesaw framework, where a discrete symmetry structure gives rise to a distinctive neutrino mass texture. This texture imposes strong constraints on the neutrino mixing parameters. Moreover, the model naturally reproduces the observed charged lepton mass hierarchy through symmetry protected suppressions. We also explore the phenomenological consequences of the texture, particularly its imprints on low energy observables such as neutrinoless double beta($0\nu\beta\beta$) decay and charged lepton flavour violation (cLFV).
The structure of the paper is as follows: in Section II, we introduce the model and the underlying symmetry groups, and discuss the construction of the model in detail. Section III contains the numerical analysis and the predictions of the model. In Sections IV and V, we discuss the implications for $0\nu\beta\beta$ and cLFV respectively. Finally, in Section VI, we present a summary and discussion of our findings.
II Theoretical Framework
We extend the Standard Model (SM) gauge group by incorporating an $A_{4}$ flavour symmetry[24, 25, 26, 27] along with cyclic $Z_{N}$[28, 29, 30, 31, 32, 33] symmetries, particularly $Z_{3}$[30, 31, 32] and $Z_{10}$[30, 31, 32]. The complete symmetry group of the model is given by
$\mathcal{G}=A_{4}\times Z_{3}\times Z_{10}\times SU(2)_{L}\times U(1)_{Y}.$ (1)
The SM field content is augmented by introducing $SU(2)_{L}$ singlet right handed neutrino fields ($\nu_{l_{R}}$) and five $SU(2)_{L}$ singlet scalar fields ($\chi,\kappa,\sigma,\rho,\zeta$). In addition, we introduce an extra $SU(2)_{L}$ doublet $\Phi$ and a triplet $\Delta$. The cyclic symmetries play a crucial role in restricting potential next to leading order (NLO) corrections, which could otherwise perturb the Yukawa Lagrangian ($\mathcal{L}_{Y}$). The transformation properties of the field content under the full symmetry of the model are summarized in Table,(1).
(a) Type-I
(b) Type-II
Figure 1: Feynman diagrams for Type-I and Type-II seesaw mechanisms.
Field $D_{l_{L}}$ $l_{R}$ $H$ $\nu_{l_{R}}$ $\Phi$ $\Delta$ $\chi$ $\kappa$ $\sigma$ $\rho$ $\zeta$
$A_{4}$ 3 (1, 1′′, 1) 3 (1, 1′′, 1) 3 3 1 1′′ 1 1 1
$Z_{3}$ $\omega$ $\omega$ 1 (1, $\omega^{*}$, $\omega^{*}$) $\omega^{*}$ $\omega^{*}$ $\omega^{*}$ $\omega^{*}$ $\omega$ $\omega$ $\omega^{*}$
$Z_{10}$ 0 (1, 4, 7) 0 (0, 2, 7) 0 0 6 6 8 3 1
$SU(2)_{L}$ 2 1 2 1 2 3 1 1 1 1 1
$U(1)_{Y}$1 -2 1 0 -1 -2 0 0 0 0 0
Table 1: Transformation properties of the fields under $\mathcal{G}$.
The leading order (LO) $\mathcal{L}_{Y}$ is constructed in the Altarelli–Feruglio basis of $A_{4}$, and is presented as shown below
$-\mathcal{L}_{Y}$ $=$ $\mathcal{L}_{l}+\mathcal{L}_{\nu}+h.c.$ (2)
Here, $\mathcal{L}_{l}$ and $\mathcal{L}_{\nu}$ correspond to the charged lepton and neutrino sectors, respectively, and are structured as follows,
$\mathcal{L}_{l}=y_{e}\,\bar{D}_{l_{L}}\,H\,e_{R}\left(\frac{\zeta}{\Lambda}\right)^{9}+\,y_{\mu}\,\bar{D}_{l_{L}}\,H\,\mu_{R}\left(\frac{\zeta}{\Lambda}\right)^{6}+\,y_{\tau}\,\bar{D}_{l_{L}}\,H\,\tau_{R}\left(\frac{\zeta}{\Lambda}\right)^{3},$ (3)
and
$\mathcal{L}_{\nu}$ $=$ $y_{D_{1}}\bar{D}_{l_{L}}\Phi\,\nu_{e_{R}}+\,y_{D_{2}}\bar{D}_{l_{L}}\Phi\,\nu_{\mu_{R}}\left(\frac{\sigma}{\Lambda}\right)+\,y_{D_{3}}\bar{D}_{l_{L}}\Phi\,\nu_{\tau_{R}}\left(\frac{\rho}{\Lambda}\right)+\frac{1}{2}\overline{\nu^{C}_{e_{R}}}\nu_{e_{R}}+\frac{1}{2}y_{1}\overline{\nu^{C}_{\mu_{R}}}\nu_{\tau_{R}}\zeta+\frac{1}{2}y_{1}\overline{\nu^{C}_{\tau_{R}}}\nu_{\mu_{R}}\zeta+$ (4)
$\frac{1}{2}y_{2}\overline{\nu^{C}_{\mu_{R}}}\nu_{\mu_{R}}\kappa+\frac{1}{2}y_{3}\overline{\nu^{C}_{\tau_{R}}}\nu_{\tau_{R}}\chi+y_{\Delta}\bar{D}_{l_{L}}D^{C}_{l_{L}}i\sigma_{2}\Delta,$
where $\Lambda$ is the effective scale of the theory. The symmetry of $\mathcal{L}_{Y}$ is spontaneously broken when the scalar fields acquire their respective vevs. A possible complete symmetry breaking pattern of the present model is shown below
$\begin{array}[]{c}\mathcal{G}\\[7.0pt] \Big{\downarrow}\ \langle\zeta\rangle\\[7.0pt] SU(2)_{L}\times U(1)_{Y}\times A_{4}\\[7.0pt] \Big{\downarrow}\ \langle\sigma,\,\rho,\,\chi,\,\kappa\rangle\\[7.0pt] SU(2)_{L}\times U(1)_{Y}\\[7.0pt] \Big{\downarrow}\ \langle\Phi,\,\Delta\rangle\langle H\rangle\\[7.0pt] U(1)_{\text{em}}\end{array}$ (5)
It is to be noted that the $Z_{3}$ and $Z_{10}$ symmetries of $\mathcal{L}_{Y}$ is broken by the vev of the scalar singlet $\zeta$. This breaking is important for the explanation of the charged lepton mass hierarchy  [34, 35]. For the following vacuum alignments: $\langle H\rangle=v_{h}(1,0,0)^{T}$, $\langle\Phi\rangle=v_{\Phi}(0,0,1)^{T}$, $\langle\Delta\rangle=v_{\Delta}(-1,0,-1)^{T}$, $\langle\chi\rangle=v_{\chi}$, $\langle\kappa\rangle=v_{\kappa}$, $\langle\sigma\rangle=v_{\sigma}$, $\langle\rho\rangle=v_{\rho}$, $\langle\zeta\rangle=v_{\zeta}$, one can derive the structure of the fermion mass matrices. The model predicts a diagonal charged lepton mass matrix, with masses satisfying the hierarchical relation
$m_{e}:m_{\mu}:m_{\tau}=\lambda_{f}^{6}\,y_{e}:\lambda_{f}^{3}\,y_{\mu}:y_{\tau},$ (6)
where $\lambda_{f}=v_{\zeta}/\Lambda$. This relation can successfully account for the observed hierarchy [36, 37] among the charged lepton masses.
The complex symmetric Majorana neutrino mass matrix in the Type-I+II seesaw framework is given by $M_{\nu}=M_{II}-M_{D}M_{R}^{-1}M_{D}^{T}$, where $M_{D}$, $M_{R}$, and $M_{II}$ denote the Dirac neutrino mass matrix, the right handed Majorana neutrino mass matrix, and the mass matrix arising from the Type-II seesaw mechanism, respectively. In the present model, the effective Type-I+II neutrino mass matrix takes the following form
$M_{\nu}=\begin{pmatrix}r-2t&t&-q\\ t&-p&t\\ -q&t&s-2t\end{pmatrix},$ (7)
where
$p$ $=$ $\frac{y^{2}_{D_{1}}v^{2}_{\Phi}}{M_{1}},$ (8)
$q$ $=$ $\frac{y_{D_{2}}y_{D_{3}}y_{1}v^{2}_{\Phi}v_{\rho}v_{\sigma}v_{\zeta}}{\left((y_{1}v_{\zeta})^{2}-y_{2}y_{3}v_{\kappa}v_{\chi}\right)\Lambda^{2}},$ (9)
$r$ $=$ $\frac{y^{2}_{D_{3}}y_{2}v^{2}_{\Phi}v^{2}_{\rho}v_{\kappa}}{\left((y_{1}v_{\zeta})^{2}-y_{2}y_{3}v_{\kappa}v_{\chi}\right)\Lambda^{2}},$ (10)
$s$ $=$ $\frac{y^{2}_{D_{2}}y_{3}v^{2}_{\Phi}v^{2}_{\sigma}v_{\chi}}{\left((y_{1}v_{\zeta})^{2}-y_{2}y_{3}v_{\kappa}v_{\chi}\right)\Lambda^{2}}.$ (11)
For the sake of simplicity, we refer to the parameters on the right hand side of Eqs. (8)–(11) as model parameters, and those on the left hand side as texture parameters.
The texture in Eq. (7) features a unique relation
$(M_{\nu}){e\mu}=(M{\nu})_{\mu\tau}.$ (12)
We refer to this texture as the Echo 12–23 Texture, as the (12) and (23) elements of $M_{\nu}$ echo each other through their equality. This correlation may significantly restrict the allowed values of the observable parameters. Before delving into the predictive implications of this texture, we first briefly review the parametrization of the PMNS matrix  [38]. According to the Particle Data Group (PDG) [36], the Pontecorvo Maki Nakagawa Sakata (PMNS) matrix ($U$) is parametrized in the following way
$U=P_{\phi}\,\tilde{U}\,P_{M},$ (13)
where $P_{\phi}=\mathrm{diag}(e^{i\phi_{1}},e^{i\phi_{2}},e^{i\phi_{3}})$ contains the unphysical phases, and $P_{M}=\mathrm{diag}(e^{i\alpha},e^{i\beta},1)$ encodes the Majorana phases. The arbitrary phases $\phi_{i}$ ($i=1,2,3$) are absorbed by redefining the charged lepton fields, simplifying the PMNS matrix to
$U=\tilde{U}P_{M}.$ (14)
In the charged lepton diagonal basis, $U$ diagonalizes $M_{\nu}$ as
$U^{T}M_{\nu}U=\mathrm{diag}(m_{1},m_{2},m_{3}).$ (15)
This leads to the equivalent condition
$\tilde{U}^{T}M_{\nu}\tilde{U}=\mathrm{diag}(\tilde{m}_{1},\tilde{m}_{2},m_{3}),$ (16)
where the rephased masses are defined as $\tilde{m}_{1}=m_{1}e^{-2i\alpha}$ and $\tilde{m}_{2}=m_{2}e^{-2i\beta}$. Thus, $\tilde{U}$ is the matrix that diagonalizes the $M_{\nu}$ of our model. The general form of $\tilde{U}$ is shown below
$\tilde{U}=\begin{bmatrix}c_{12}c_{13}&s_{12}c_{13}&s_{13}e^{-i\delta}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}e^{i\delta}&c_{12}c_{23}-s_{12}s_{13}s_{23}e^{i\delta}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}e^{i\delta}&-c_{12}s_{23}-s_{12}s_{13}c_{23}e^{i\delta}&c_{13}c_{23}\end{bmatrix},$ (17)
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}$. The $\theta_{ij}$ represents the three mixing angles.
The diagonalizing matrix $\tilde{U}$ imposes constraints on the neutrino mass matrix $M_{\nu}$, such that all complex texture parameters $p$, $q$, $r$, $s$, and $t$ can be expressed in terms of only three real parameters: $Re[s],Im[s]$ and $Re[t]$. In other words, these three free parameters are sufficient to fully determine $M_{\nu}$ and its associated phenomenological predictions.
III Numerical Analysis
We first perform a numerical analysis in the light of normal hierarchy (NH) of neutrino masses. In this regard, the $3\sigma$ ranges of the mixing angles and the Dirac CP phase $\delta$ are used as inputs. A large sample of points in the parameter space ${Re[s],Im[s],Re[t]}$ is generated, requiring that the mass-squared differences $\Delta m_{21}^{2}$ and $\Delta m_{31}^{2}$ lie within the $3\sigma$ ranges reported by NuFIT 6.0 [39]. In addition, the sum of the three light neutrino masses is required to satisfy the cosmological bound $\sum m_{i}<0.12\,\text{eV}$[40]. From the dataset, we find that viable parameter points exist within the following ranges
$(Re[s])_{\text{min}}=0.0752\,\text{eV},\quad(Re[s])_{\text{max}}=0.0843\,\text{eV};$
$(Im[s])_{\text{min}}=-0.0089\,\text{eV},\quad(Im[s])_{\text{max}}=0.0086\,\text{eV};$
$(Re[t])_{\text{min}}=0.0192\,\text{eV},\quad(Re[t])_{\text{max}}=0.0236\,\text{eV},$
However, when performing a similar scan assuming an inverted hierarchy (IH), we find that the model yields no viable parameter points. Thus, IH is disfavoured in this model.
It is important to note that, although the mixing angles and $\delta$ are scanned over their full $3\sigma$ ranges, we observe that for the viable parameter space yielding $\sum m_{i}$ within $[0.1,0.118]$ eV, the Dirac CP phase $\delta$ is strongly restricted. A large portion of its allowed range, approximately between $104^{\circ}$ and $285.6^{\circ}$, is forbidden by the model. While $\delta$ is typically treated as a free parameter in many flavour models, here its viable values are tightly limited due to the structure of the Echo 12–23 Texture. As a result, the overall parameter space through this restriction on $\delta$ is also constrained. From the predicted spectra of the neutrino mass eigenvalues, $m_{1}$ and $m_{2}$ appears to overlap in certain regions, potentially giving the impression of an exact degeneracy. However, when we calculate $m_{2}-m_{1}$, the difference is never zero, thereby removing the notion of exact degeneracy. This supports a quasi-degenerate mass spectra, i.e ., $m_{1}\lesssim m_{2}\lesssim m_{3}$. The model predicts the Majorana phases within a narrow range. The phases $\alpha$ and $\beta$ are constrained to lie between $-90^{\circ}$ and $90^{\circ}$ and $-12.7^{\circ}$ and $13.6^{\circ}$, respectively. Since the Dirac CP phase $\delta$ is subject to restrictions, we also study the associated Jarlskog invariant $J_{\text{CP}}$, which quantifies leptonic CP violation and depends directly on $\delta$ and the mixing angles.
Figure 2: The plots of the predicted parameters are shown. (a) $\delta$ vs $\sum m_{i}$. (b) $m_{2,1}$ vs $m_{3}$. (c) $m_{2}-m_{1}$ vs $m_{1}$. (d) $\alpha$ vs $\beta$. (e) $\delta$ vs $J_{\text{CP}}$.
The $J_{\text{CP}}$ is predicted to lie between $-0.0324$ and $0.030$. To summarize our findings, we list the numerical values of the predicted parameters in Table 2. To have a graphical understanding of the predicted parameters, we highlight the plots for the same in Fig. 2. In Fig. 3, we highlight the plots of parameters, $Re[s],Im[s],Re[t]$.
Prediction $\delta$ $m_{1}$ $m_{2}$ $m_{3}$ $\sum m_{i}$ $\alpha$ $\beta$ $J_{\text{CP}}$
Minimum $285.7^{\circ}$ 0.021 eV 0.023 eV 0.053 eV 0.1 eV $-90^{\circ}$ $-12.7^{\circ}$ $-0.0324$
Maximum $421^{\circ}$ 0.030 eV 0.031 eV 0.058 eV 0.118 eV $90^{\circ}$ $13.6^{\circ}$ $0.030$
Table 2: The approximate minimum and maximum values of $\delta$, $m_{1}$, $m_{2}$, $m_{3}$, $\sum m_{i}$, $\alpha$, $\beta$, and $J_{\text{CP}}$ as predicted by the model.
So far, we have presented the predictions for the low energy observable quantities as well as the texture parameters $Re[s]$, $Im[s]$, and $Re[t]$ that define the entire neutrino mass matrix and all the predictions associated with it. These texture parameters, however, are not arbitrary, they are generated from a deeper set of model parameters, including the Yukawa couplings and flavon vevs, as outlined in Eqs. (8)–(11). Therefore, by using the predicted values of the texture parameters, we can determine or constrain the underlying model parameters that define the theory at a more fundamental level. Though the framework cannot directly predict the individual model parameters, it can provide numerical information for several interesting combinations of them.
Figure 3: The plots of the free parameters are shown. (a)$Im[s]$ vs $Re[s]$. (b) $Im[s]$ vs $Re[t]$.
In this connection, we highlight the model parameter combinations below, which are functions of the free parameters $Re[s]$, $Im[s]$, and $Re[t]$
$G_{1}=\frac{y_{D_{1}}v_{\Phi}}{\sqrt{M_{1}}},\quad G_{2}=\frac{y_{D_{2}}v_{\Phi}v_{\sigma}}{\sqrt{y_{2}v_{\kappa}}\Lambda},\quad G_{3}=\frac{y_{D_{3}}v_{\Phi}v_{\rho}}{\sqrt{y_{3}v_{\chi}}\Lambda},$
$\hskip 42.67912ptG_{4}=\frac{y_{1}v_{\zeta}}{\sqrt{y_{2}y_{3}v_{\kappa}v_{\chi}}},\quad G_{5}=y_{\Delta}v_{\Delta}.$ (18)
The predicted numerical values of the model parameter combinations $G_{1},G_{2},G_{3},G_{4}$, and $G_{5}$ are listed in Table 3.
Parameter $|G_{1}|$ $|G_{2}|$ $|G_{3}|$ $|G_{4}|$ $|G_{5}|$ $Arg[G_{1}]$ $Arg[G_{2}]$ $Arg[G_{3}]$ $Arg[G_{4}]$ $Arg[G_{5}]$
Minimum 0.156 $\text{eV}^{\frac{1}{2}}$ 0.05 $\text{eV}^{\frac{1}{2}}$ 0.025 $\text{eV}^{\frac{1}{2}}$ 0.118 0.060 eV $-90.000^{\circ}$ $-92.364^{\circ}$ $-92.854^{\circ}$ $-42.100^{\circ}$ $-21.170^{\circ}$
Maximum 0.253 $\text{eV}^{\frac{1}{2}}$ 0.355 $\text{eV}^{\frac{1}{2}}$ 0.224 $\text{eV}^{\frac{1}{2}}$ 1.185 0.0722 eV $90.000^{\circ}$ $92.233^{\circ}$ $92.767^{\circ}$ $50.000^{\circ}$ $23.000^{\circ}$
Table 3: The approximate maximum and minimum values of the model parameter combinations.
To visualise these combinations graphically, we plot them in Fig. 4.
Figure 4: The plots of the model parameter combinations are shown.
The model is now equipped with all the necessary numerical inputs to examine its consistency in light of low energy phenomenologies such as $0\nu\beta\beta$ decay and cLFV.
IV Neutrinoless Double Beta Decay
The model adheres to the Majorana nature of the neutrinos, therefore, it is pertinent to discuss the effective Majorana mass, $m_{\beta\beta}$. This quantity is a crucial observable in determining the half life ($T_{1/2}$) of $0\nu\beta\beta$ decay [41, 42, 43]. The $0\nu\beta\beta$ decay is a hypothetical decay in which an unstable nucleus undergoes two simultaneous beta decays without emitting neutrinos. Such a decay is only possible if neutrinos are Majorana particles. A typical Feynman diagram for this process is shown in Fig. 5. A confirmed observation of this process would thus provide direct evidence for the Majorana nature of neutrinos. Although no positive signal has been observed to date, several experiments have placed stringent upper limits on the observable $m_{\beta\beta}$, which are summarised in Table 4. The general expression for $m_{\beta\beta}$ is given below
$m_{\beta\beta}=|U^{2}_{ei}m_{i}|,\quad\text{where,}\,\,i=1,2,3.$ (19)
Figure 5: Feynman diagram for $0\nu\beta\beta$ decay.
Isotope ${}^{130}\mathrm{Te}$ ${}^{136}\mathrm{Xe}$ ${}^{76}\mathrm{Ge}$ ${}^{136}\mathrm{Xe}$ ${}^{76}\mathrm{Ge}$
[…]
Collaboration CUORE [44] EXO [45] GERDA [46] KamLAND-Zen [47, 48] LEGEND-1000 [49]
Table 4: The present upper limits on $m_{\beta\beta}$ from $0\nu\beta\beta$ decay searches in different isotopes.
Figure 6: The plot of $m_{\beta\beta}$ versus $m_{1}$. In the plot, the ‘yellow’ area show the generic region allowed by current oscillation data in case of NH. The ‘gray’ band indicates the bounds on $m_{\beta\beta}$ set by KamLAND-Zen and GERDA. The ‘faded pink’ band corresponds to the region excluded by cosmology in case of NH. The ‘light green’ band shows the projected sensitivity of the LEGEND-1000, and the ‘red’ curve represents the prediction of the present model.
As mentioned previously, $Re[s]$, $Im[s]$, and $Re[t]$ dictate the structure of $M_{\nu}$ and all associated predictions. Hence, $m_{\beta\beta}$ is no exception. For NH, the model places stringent bound on $m_{\beta\beta}$, predicting it to lie within the range 0.0060 eV – 0.0227 eV. This range is consistent with the current upper limits set by KamLAND-Zen and GERDA, and lies within the projected sensitivity of the LEGEND-1000 experiment. To visualise the prediction of $m_{\beta\beta}$ by the model, we show the plot of $m_{\beta\beta}$ versus the lightest neutrino mass eigenvalue ($m_{1}$) in Fig. 6
V Charged Lepton Flavour Violation
In the SM of particle physics, cLFV processes [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63] are highly suppressed, with predicted branching ratios around $10^{-50}$. Therefore, cLFV is one of the most promising probes of physics beyond the SM. Among the various decay channels, transitions involving muons are the most interesting, as muons have a longer lifetime than other leptons and are abundantly produced in cosmic radiation. This makes muon decays ideal for study. So far, there is no experimental evidence of cLFV decays. However, ongoing experiments are steadily improving their sensitivities to search for these rare processes. To complement these experimental efforts, it is worthwhile to explore theoretical frameworks that could explain or predict such decays. As an extension of our model, we plan to study the dominant decay $\mu\rightarrow e\gamma$. The current best limit on the branching ratio of the $\mu\rightarrow e\gamma$ decay is provided by MEG [64]
$\text{BR}(\mu\rightarrow e\gamma)<4.2\times 10^{-13}.$ (20)
The MEG II [65, 66, 67] experiment is currently searching for this rare decay at the Paul Scherrer Institute (PSI) muon beam facility. Its goal is to improve the sensitivity of measurements to $BR(\mu\rightarrow e\gamma)\sim 6\times 10^{-14}$ by the end of 2026. In the present model, it is possible to explore how each seesaw mechanism may contribute to the $\mu\rightarrow e\gamma$ decay. We then aim to determine which contribution is the dominant one. In the case of Type-I seesaw, the $\mu\rightarrow e\gamma$ decay arises due to the existence of heavy-light neutrino mixing. To understand this, we write the effective Lagrangian for the Type-I seesaw in the following manner
$\mathcal{L}_{\nu}$ $=$ $\frac{1}{2}\overline{n^{c}}Mn+h.c,$ (21)
with
$n$ $=$ $\begin{pmatrix}\nu^{c}_{L}\\ \nu_{R}\\ \end{pmatrix},$ (22)
and the $6\times 6$ mass matrix,
$M$ $=$ $\begin{pmatrix}0&M_{D}\\ M^{T}_{D}&M_{R}\\ \end{pmatrix}.$ (23)
The block diagonalisation of $M$ is achieved by the following transformation
$\begin{pmatrix}1-\frac{RR^{\dagger}}{2}&R\\ -R^{\dagger}&1-\frac{RR^{\dagger}}{2}\\ \end{pmatrix}^{T}\begin{pmatrix}0&M_{D}\\ M^{T}_{D}&M_{R}\\ \end{pmatrix}\begin{pmatrix}1-\frac{RR^{\dagger}}{2}&R\\ -R^{\dagger}&1-\frac{RR^{\dagger}}{2}\\ \end{pmatrix}\simeq\begin{pmatrix}m_{\text{light}}&0\\ 0&M_{\text{heavy}}\\ \end{pmatrix},$ (24)
where $R=M^{\dagger}_{D}M_{R}^{-1}$, $m_{\text{light}}=-M_{D}M_{R}^{-1}M_{D}^{T}$ and $M_{\text{heavy}}=M_{R}$. The next step is to diagonalise the block diagonalised matrix in Eq. 24. This is achieved via the transformation below
$\begin{pmatrix}U&0\\ 0&U_{R}\\ \end{pmatrix}^{T}\begin{pmatrix}m_{\text{light}}&0\\ 0&M_{\text{heavy}}\\ \end{pmatrix}\begin{pmatrix}U&0\\ 0&U_{R}\\ \end{pmatrix}\simeq\begin{pmatrix}\text{diag}(m_{i})&0\\ 0&\text{diag}(M_{i})\\ \end{pmatrix},$ (25)
where $U$ is the PMNS matrix that diagonalises $m_{\text{light}}$, and $U_{R}$ diagonalises $M_{\text{heavy}}$. This two step diagonalisation of $M$ can be simplified to a single step in the following way
$\begin{pmatrix}U&0\\ 0&U_{R}\\ \end{pmatrix}^{T}\begin{pmatrix}1-\frac{RR^{\dagger}}{2}&R\\ -R^{\dagger}&1-\frac{RR^{\dagger}}{2}\\ \end{pmatrix}^{T}\begin{pmatrix}0&M_{D}\\ M^{T}_{D}&M_{R}\\ \end{pmatrix}\begin{pmatrix}1-\frac{RR^{\dagger}}{2}&R\\ -R^{\dagger}&1-\frac{RR^{\dagger}}{2}\\ \end{pmatrix}\begin{pmatrix}U&0\\ 0&U_{R}\\ \end{pmatrix}\simeq\begin{pmatrix}\text{diag}(m_{i})&0\\ 0&\text{diag}(M_{i})\\ \end{pmatrix}.$ (26)
Therefore, $M$ can be diagonalised by the matrix
$D=\begin{pmatrix}(1-\frac{RR^{\dagger}}{2})U&RU_{R}\\ -R^{\dagger}U&(1-\frac{RR^{\dagger}}{2})U_{R}\\ \end{pmatrix},$ (27)
and the off diagonal part $K=R\,U_{R}=M^{\dagger}_{D}M_{R}^{-1}U_{R}$ carries the information of heavy-light neutrino mixing. In the case of the Type-I seesaw, $K$ directly enters the expression for the branching ratio of the $\mu\rightarrow e\gamma$ decay, which is given by [68, 69, 70, 71, 72, 73]
$BR(\mu\rightarrow e\gamma)=\frac{3\alpha_{f}}{8\pi}\left|\sum_{i}K_{ei}K^{\dagger}_{i\mu}F\left(\frac{M_{i}^{2}}{M_{W}^{2}}\right)\right|^{2},$ (28)
where, $\alpha_{f}$ stands for the fine structure constant, $K$ corresponds to the heavy-light mixing matrix, $M_{i}$ is the mass of the right handed neutrinos and $M_{W}$ is the mass of $W$ boson. The factor $F(x)$ is expressed as
$F(x)=\frac{x(1-6x+3x^{2}+2x^{3}-6x^{2}\ln{x})}{2(1-x)^{4}},$ (29)
where, $x=\frac{M_{i}^{2}}{M_{W}^{2}}$. To compute $BR(\mu\rightarrow e\gamma)$, we consider $M_{R}$ in the TeV range, requiring $M_{D}\sim$ MeV to reproduce light neutrino masses $\sim 10^{-2}$ eV. This leads to $BR(\mu\rightarrow e\gamma)\sim\mathcal{O}(10^{-40})$, far below the experimental limit. Hence, the Type-I contribution to $BR(\mu\rightarrow e\gamma)$ is insignificant.
With this inference in mind, we now turn our attention to the Type-II seesaw contribution to $BR(\mu\rightarrow e\gamma)$. In this case, the process arises at the one loop level, mediated by the singly and doubly charged components of the scalar triplet $\Delta$. The corresponding branching ratio is given by [74, 63, 75]
$BR(\mu\rightarrow e\gamma)=\frac{\alpha_{f}\left|(Y^{*}_{II}Y_{II})_{e\mu}\right|^{2}}{192\pi G_{F}^{2}}\left(\frac{1}{m^{2}_{\Delta^{\pm}}}+\frac{8}{m^{2}_{\Delta^{\pm\pm}}}\right)^{2},$ (30)
where $Y_{II}$ is the Type-II Yukawa matrix, and $G_{F}$ denotes the Fermi constant. In our analysis, we assume that the charged components of the scalar triplet are approximately degenerate in mass, i.e., $m_{\Delta^{++}}\simeq m_{\Delta^{+}}\simeq m_{\Delta}$. This assumption is well justified in scenarios where the mass splitting among triplet components is small. Substituting the model predicted value of $(Y^{*}_{II}Y_{II})_{e\mu}$ into Eq. (30), the branching ratio simplifies to
$BR(\mu\rightarrow e\gamma)\simeq 0.0208\,\frac{\alpha_{f}}{\pi G_{F}^{2}}\left(\frac{\left|G_{5}\right|}{m_{\Delta}v_{\Delta}}\right)^{4}.$ (31)
From Eq. (31), it is evident that the branching ratio depends on the parameters $|G_{5}|$, $m_{\Delta}$, and $v_{\Delta}$. Using the numerical values of $|G_{5}|$ obtained from our model, we perform a scan over $m_{\Delta}$ and $v_{\Delta}$ to identify the regions of parameter space that yield branching ratios consistent with experimental limits. This allows us to determine the most favourable regions in the $(m_{\Delta},v_{\Delta},|G_{5}|)$ parameter space. Based on our numerical analysis, we find that the predicted branching ratio $BR(\mu\rightarrow e\gamma)$ lies within the range $[10^{-15},4\times 10^{-13}]$, which is compatible with the current MEG bound and the projected sensitivity of MEG-II. While $|G_{5}|$ remains fixed within its allowed range, $[0.06,0.072]$, three distinct and viable sets of bounds for $m_{\Delta}$ and $v_{\Delta}$ emerge, as summarised in Table 5.
           $|G_{5}|\in[0.06,0.072]$   $BR(\mu\rightarrow e\gamma)\in[10^{-15},4\times 10^{-13}]$
           $m_{\Delta}$            $v_{\Delta}$
           Set            Minimum            Maximum            Minimum            Maximum
           A            870 GeV            4 TeV            2.30 eV            2.50 eV
           B            1.00 TeV            4.50 TeV            1.98 eV            2.10 eV
           C            1.20 TeV            8.70 TeV            1.00 eV            1.70 eV
Table 5: Numerical sets of $(m_{\Delta},v_{\Delta})$ yielding $BR(\mu\rightarrow e\gamma)\in[10^{-15},4\times 10^{-13}]$ for $|G_{5}|\in[0.06,0.072]$.
Figure 7: The plot of $BR(\mu\rightarrow e\gamma)$ vs. $m_{\Delta}$. The ‘red’, ‘blue’, and ‘faded blue’ points represent the predicted ranges of $BR(\mu\rightarrow e\gamma)$ for set A, set B, and set C, respectively. The ‘faded red’ region is excluded by the MEG experiment, while the ‘pale green’ region indicates the $BR(\mu\rightarrow e\gamma)$ values that are within reach of MEG-II sensitivity. The potential reach of high-energy colliders such as HL-LHC, HE-LHC, and FCC-hh is also overlaid for comparison.
The HL-LHC [76], HE-LHC [77] and FCC-hh [78] are expected to probe the higher mass range of the scalar triplet $\Delta$. In particular, the collider signatures of $\Delta^{\pm\pm}$[79], such as same-sign dilepton decays ($\Delta^{\pm\pm}\to\ell^{\pm}\ell^{\pm}$) for small $v_{\Delta}$, and diboson decays ($\Delta^{\pm\pm}\to W^{\pm}W^{\pm}$) for larger $v_{\Delta}$, provide promising avenues to test the scalar triplet mass. These searches, combined with charged lepton flavour violation, can play a crucial role in probing the parameter space of our model. To visualize our findings, we present a plot of $BR(\mu\rightarrow e\gamma)$ as a function of $m_{\Delta}$ in Fig. 7. Therefore, we may infer that in our model the dominant contribution to the $BR(\mu\rightarrow e\gamma)$ comes from the Type-II seesaw mechanism.
VI Summary and Discussions
We have presented a flavour driven neutrino mass model that stands out by generating a distinctive and predictive Echo 12–23 Texture, realized from first principles within a Type-I+II seesaw framework. This elegant texture emerges naturally from an extended symmetry structure, $A_{4}\times Z_{3}\times Z_{10}$, which plays a crucial role in shaping the Yukawa sector and suppressing unwanted contributions via higher dimensional operators. Remarkably, the model accounts for the lepton mass hierarchy and predicts the entire neutrino mass matrix using just three real input parameters. The model makes sharp predictions for the three neutrino mass eigenvalues, as well as the Dirac and Majorana CP phases. Furthermore, it yields consistent predictions for the effective neutrino mass in light of $0\nu\beta\beta$ decay and the branching ratio of the dominant $\mu\rightarrow e\gamma$ decay. This opens up prospects for verification in current and upcoming experiments. Overall, this framework offers a fresh path toward decoding the flavour puzzle in the lepton sector.
Acknowledgement
The research work of MD is financially supported by Council of Scientific and Industrial Research (CSIR), Government of India through a NET Senior Research Fellowship vide grant No. 09/0059(15346)/2022-EMR-I.